OpenStudy (anonymous):

What is the correct reason for statement 2? Prove: 10(w + 6) + 4 = 10w + 64 Statements Reasons 1. 10(w + 6) + 4 1. Given 2. (10w + 60) + 4 2. 3. 10w + (60 + 4) 3. Associative Property 4. 10w + 64 4. Addition A. Distributive Property B. Commutative Property C. Associative Property of Multiplication D. Addition

OpenStudy (anonymous):

@jessiehuang_

OpenStudy (calculusxy):

Distributive property.

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

is he right @jessiehuang_

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

What is the correct reason for statement 2? Prove: 1 + 3x + (x ∙ 4) = 7x + 1 Statements Reasons 1. 1 + 3x + (x ∙ 4) 1. Given 2. 1 + 3x + 4x 2. 3. 1 + 7x 3. Addition 4. 7x + 1 4. Commutative Property A. Associative Property B. Multiplication C. Commutative Property D. Addition

OpenStudy (anonymous):

b multiplication

OpenStudy (anonymous):

ty i have another one do u like helping me??

OpenStudy (anonymous):

I don't mind are you in alg 1 or geometry?

OpenStudy (anonymous):

alg 1

OpenStudy (anonymous):

What is the correct reason for statement 3? Prove: (11m) ∙ 4 + m = 45m Statements Reasons 1. (11m) ∙ 4 + m 1. Given 2. 4 ∙ (11m) + m 2. Commutative Property 3. (4 ∙ 11) ∙ m + m 3. 4. 44m + m 4. Commutative Property 5. 45m 5. Addition A. Multiplication B. Distributive Property C. Associative Property D. Addition

OpenStudy (anonymous):

c

OpenStudy (anonymous):

What is the correct reason for statement 5? Prove: k + 4 + 6(1 + k) = 11k + 10 Statements Reasons 1. k + 4 + 6(1 + k) 1. Given 2. k + 4 + (6+ 6k) 2. Distributive Property 3. k + (4 + 6) + 6k 3. Associative Property 4. k + 10 + 6k 4. Addition 5. k + 6k + 10 5. 6. 7k + 10 6. Addition A. Multiplication B. Commutative Property C. Associative Property D. Addition

OpenStudy (anonymous):

b commutative

OpenStudy (anonymous):

Select the correct reasons for statements 1 and 3 to complete the proof. Prove: 4x + 8 • x • 3 = 28x 1. 4x + 8 • x • 3 = 4x + 8 • 3 • x 1. ? 2. 4x + 8 • x • 3 = 4x + 24x 2. Multiplication 3. 4x + 8 • x • 3 = 28x 3. ? A. Associative Property of Addition, Addition B. Commutative Property of Addition, Multiplication C. Associative Property of Multiplication, Addition D. Commutative Property of Multiplication, Addition

OpenStudy (anonymous):

d

OpenStudy (anonymous):

i just turned it in

OpenStudy (anonymous):

i got 80% i passed

OpenStudy (anonymous):

im bout to close this one and open another 1

OpenStudy (anonymous):

ok